Hi. I'm trying very hard to figure out the basal metabolic rate of a rat.
The rat weight is 250 grams, which I converted to .25 kg.
The ml O2 used in 1 minute was 7.2 ml.
The ml O2 used in 1 hour equals 432 ml O2/hr.
Now I am supposed to get metabolic rate by ml O2/hr over weight in kg = ml O2/kg/hr. I don't understand how to get the answer.

Answer by Alan3354(69443)

Hi. I'm trying very hard to figure out the basal metabolic rate of a rat.
The rat weight is 250 grams, which I converted to .25 kg.
The ml O2 used in 1 minute was 7.2 ml.
The ml O2 used in 1 hour equals 432 ml O2/hr.
Now I am supposed to get metabolic rate by ml O2/hr over weight in kg = ml O2/kg/hr. I don't understand how to get the answer.
=====================
ml/hr = 432
-----
ml/hr / kg = 432/0.25 = 1728